OpenStudy (anonymous):

Use the rules of derivatives to find the derivative of P(t)=t^2(t^2+1)^5/2

OpenStudy (anonymous):

product rule plus chain rule

OpenStudy (anonymous):

Great! I couldn't decide if it was chain rule first, thanks

OpenStudy (anonymous):

\[(fg)'=f'g+g'f\] with \[f(t)=t^2, f'(t)=2t, g(t)=(t^2+1)^{\frac{5}{2}}, g'(t)=\frac{5}{2}(t^2+1)^{\frac{3}{2}}\times 2t\]

OpenStudy (anonymous):

some algebra will simplify this, that is just the start